# Heights and periods of algebraic cycles in families

### v1

@@ C.2. Betti foliation
The holomorphic vector bundle ${\mathscr{V}}$ is endowed with a connection $\nabla$ whose sheaf of the horizontal sections is ${\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}}{\mathbb{C}}$.
@@
Locally on $S$, the local system ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}$ is trivial and so is $({\mathscr{V}},\nabla)|_{\Delta}$. Thus ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}\subseteq{\mathscr{V}}_{\Delta}$ becomes ${\mathbb{Z}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$, which extends to ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$. Notice that $({\mathbb{R}}^{2d}\times\Delta)\cap{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}$ is 00 on each fiber by weight reasons. Hence ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta\rightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)$ is a real analytic diffeomorphism, and over each $s\in\Delta$ it becomes a group homomorphism. Moreover, the image of $\{r\}\times\Delta$ is easily seen to be complex analytic for any $r\in{\mathbb{R}}^{2d}$.
@@
This gives a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which we call the Betti foliation. More concretely ${\mathscr{F}}_{\mathrm{Betti}}$ is defined as follows: for each $x\in\pi^{-1}(\Delta)$, the local leaf through $x$ is the fiber $b_{\Delta}^{-1}(b_{\Delta}(x))$. Each leaf is holomorphic to the discussion above. In fact, the Betti foliation is the unique foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which is everywhere transverse to the fibers of $\pi$ and whose set of leaves contains all torsion multisections.
For each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, the Betti foliation induces a decomposition $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$, and $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ induces a natural isomorphism $T_{x}{\mathscr{F}}_{\mathrm{Betti}}\cong T_{\pi(x)}S$. The translation on the torus ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ yields a canonical isomorphism $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})$. Hence, we have a linear map
@@ C.5. Normal functions arising from families of algebraic cycles
Now, we turn to the family version and define the corresponding normal function. Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$. Let $Z$ be a family of homologically trivial $n$-cocycle in $X/S$, i.e. a formal sum of integral subschemes of $X$ which are flat and dominant over $S$ such that each fiber $Z_{s}$ is a homologically trivial $n$-cocycle of $X_{s}$.
@@
$\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{n}(X/S),\qquad s\mapsto{\mathrm{AJ}}(Z_{s})$
It is an admissible normal function.
The following proposition is a simple application of the Abel–Jacobi map and the Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on intermediate Jacobians introduced in §$\S$§C.2.
@@
Set $S^{\circ}:=S\setminus S_{{\mathscr{F}}}(1)$.
Assume $X/S$ and $Z$ are defined over $\overline{\mathbb{Q}}$. Then $[Z_{s}]\in\mathrm{CH}^{n}(X_{s})$ is non-torsion for any $s\in S^{\circ}({\mathbb{C}})\setminus S^{\circ}(\overline{\mathbb{Q}})$.
@@
Assume $[Z_{s}]$ is a torsion point of $\mathrm{CH}^{n}(X_{s})$. Then $[Z_{t}]$ is a torsion point of $\mathrm{CH}^{n}(X_{t})$ for each $t\in\bar{s}(\overline{\mathbb{Q}})$, and $t\mapsto{\mathrm{AJ}}(Z_{t})$ is a torsion section of ${\mathscr{J}}^{n}(X/S)\times_{S}\bar{s}\rightarrow\bar{s}$. So $\nu_{Z}(\bar{s})\subseteq{\mathscr{F}}_{\mathrm{Betti}}$. If $s\not\in S(\overline{\mathbb{Q}})$, then $\dim\bar{s}\geq 1$, and hence $\bar{s}\subseteq S_{{\mathscr{F}}}(1)$ by definition. Now we are done. ∎
@@ C.6. Normal functions associated with Gross–Schoen cycles and Ceresa cycles
Let $C$ be an irreducible smooth projective curve defined over a field $k$ of genus $g\geq 3$. Let $\xi\in C(k)$. We have:
    For each subset $T\subseteq\{1,2,3\}$, the modified diagonal $\Delta_{T}(C)=\{(x_{1},x_{2},x_{3}):x_{i}=\xi\text{ for }i\not\in T,~{}x_{j}=x_{j^{\prime}}\text{ for all }j,j^{\prime}\in T\}\in\mathrm{CH}^{2}(C^{3})$.
    The (classical) Abel–Jacobi map $i_{\xi}\colon C\rightarrow\mathrm{Jac}(C)$ sending $x\mapsto[x-\xi]$.
Both $\Delta_{T}(C)$ and $i_{\xi}$ can be extended to any $\xi=\sum_{i}n_{i}e_{i}\in\mathrm{Div}^{1}(C)(k)=\mathrm{Pic}^{1}(C)(k)$. This is classical and direct for $i_{\xi}$, and for $\Delta_{T}(C)$ one can define $\Delta_{123}(C)=\{(c,c,c):c\in C\}$, $\Delta_{12}(C)=\sum n_{i}\{(c,c,e_{i}):c\in C\}$, $\Delta_{1}(C)=\sum_{i,j}n_{i}n_{j}\{(c,e_{i},e_{j}):c\in C\}$, etc.
For any $\xi\in\mathrm{Pic}^{1}(C)$, define the Gross–Schoen cycle based at $\xi$ (resp. the Ceresa cycle based at $\xi$) to be:
  • (Gross–Schoen cycle based at $\xi$) $\Delta_{{\mathrm{GS}},\xi}(C):=\Delta_{123}(C)-\Delta_{12}(C)-\Delta_{13}(C)-\Delta_{23}(C)+\Delta_{1}(C)+\Delta_{2}(C)+\Delta_{3}(C)\in\mathrm{CH}^{2}(C^{3})$.
  • (Ceresa cycle based at $\xi$) ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in\mathrm{CH}^{g-1}(\mathrm{Jac}(C))$.
${\mathrm{Ce}}_{\xi}(C)$ is clearly homologically trivial because $[-1]$ acts trivially on even-degree cohomology groups. It is not hard to check that $\Delta_{{\mathrm{GS}},\xi}(C)$ is also homologically trivial; this follows for example from [GS95, Prop. 3.1] and because the map $\mathrm{Pic}^{1}(C)\rightarrow\mathrm{CH}^{2}(C^{3})$, $\xi\mapsto\Delta_{{\mathrm{GS}},\xi}(C)$ is a group homomorphism.
Let $\omega_{C}$ be the canonical divisor on $C$.
@@
There are finitely many $\xi$’s with $(2g-2)\xi=\omega_{C}$ and each two differ from a $(2g-2)$-torsion. Hence, the Gross–Schoen cycle and the Ceresa cycle of $C$ are well-defined at $(2g-2)$-torsions.
We will associate normal functions with the Schoen cycles and Ceresa cycles. First we have ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}^{2}(C^{3})={\mathscr{J}}(H^{3}(C^{3},{\mathbb{Z}})(2))$ and ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}^{g-1}(\mathrm{Jac}(C))={\mathscr{J}}(H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1))$. Using the Poincaré duality on $\mathrm{Jac}(C)$ and on $C$, we get $H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1)=H_{3}(\mathrm{Jac}(C),{\mathbb{Z}})(-1)=\bigwedge^{3}H_{1}(C,{\mathbb{Z}})(-1)=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)$. So ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}(\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2))$.
We have $\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)\subseteq H^{3}(C^{3},{\mathbb{Z}})(2)$ in the following way. First, the Künneth formula gives a decomposition $H^{3}(C^{3},{\mathbb{Z}})=H^{1}(C,{\mathbb{Z}})^{\otimes 3}\bigoplus H^{1}(C,{\mathbb{Z}})(-1)^{\oplus 6}$. Next, $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$, as a subspace of $H^{3}(C^{3},{\mathbb{Z}})$, has a basis consisting of $\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}$, with $\alpha_{j}$ the pullback of an element in $H^{1}(C,{\mathbb{Z}})$ under the $j$-th projection $C^{3}\rightarrow C$. The symmetric group $S_{3}$ acts naturally on $C^{3}$, and this induces an action of $S_{3}$ on $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$ with $\sigma(\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3})=\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$ for each $\sigma\in S_{3}$. Then $(H^{1}(C,{\mathbb{Z}})^{\otimes 3})^{S_{3}}=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})$ for this action, with each member having a basis consisting of $\sum_{\sigma\in S_{3}}\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$.
@@
For appropriate numbers $a_{j}$. The multiplication $[n]\colon\mathrm{Jac}(C)\rightarrow\mathrm{Jac}(C)$ induces $\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})\rightarrow\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})$, $x\mapsto n^{3}x$. Hence $n^{2+j}{\mathrm{AJ}}(C_{j})=n^{3}{\mathrm{AJ}}(C_{j})$ for all $n\in{\mathbb{Z}}$, and so ${\mathrm{AJ}}(C_{j})=0$ for all $j\geq 2$.
@@
Now, we turn to defining and studying the normal functions. Let ${\mathscr{M}}_{g}$ be the moduli space of curves of genus $g$, and let $f\colon{\mathscr{C}}_{g}\rightarrow{\mathscr{M}}_{g}$ be the universal curve. We have a VHS $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}\rightarrow{\mathscr{M}}_{g}$. The fiberwise polarization $q_{H}\colon H^{1}(C,{\mathbb{Z}})\otimes H^{1}(C,{\mathbb{Z}})\xrightarrow{\cup}H^{2}(C,{\mathbb{Z}})\cong{\mathbb{Z}}(-1)$, with $\cup$ the cup product (the dual of the intersection pairing) on $H^{1}(C,{\mathbb{Z}})$, induces a polarization on $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}$.

### v2

@@ C.2. Betti foliation
Locally on $S$, the local system ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}$ is trivial and so is $({\mathscr{V}},\nabla)|_{\Delta}$. Thus ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}\subseteq{\mathscr{V}}_{\Delta}$ becomes ${\mathbb{Z}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$, which extends to ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$. Notice that $({\mathbb{R}}^{2d}\times\Delta)\cap{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}$ is 00 on each fiber by weight reasons. Hence
${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta\rightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)$
is a real analytic diffeomorphism, and over each $s\in\Delta$ it becomes a group homomorphism. Moreover, the image of $\{r\}\times\Delta$ is easily seen to be complex analytic for any $r\in{\mathbb{R}}^{2d}$.
@@
This gives a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which we call the Betti foliation. More concretely ${\mathscr{F}}_{\mathrm{Betti}}$ is defined as follows: for each $x\in\pi^{-1}(\Delta)$, the local leaf through $x$ is the fiber $b_{\Delta}^{-1}(b_{\Delta}(x))$. Each leaf is holomorphic to the discussion above. In fact, the Betti foliation is the unique foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, which is everywhere transverse to the fibers of $\pi$ and whose set of leaves contains all torsion multisections.
For each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, the Betti foliation induces a decomposition
$T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)},$
and $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ induces a natural isomorphism $T_{x}{\mathscr{F}}_{\mathrm{Betti}}\simeq T_{\pi(x)}S$. The translation on the torus ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ yields a canonical isomorphism $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})$. Hence, we have a linear map
@@ C.5. Normal functions arising from families of algebraic cycles
Now, we turn to the family version and define the corresponding normal function. Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$. Let $Z$ be a family of homologically trivial $n$-cocycles in $X/S$, i.e. a formal sum of integral subschemes of $X$ which are flat and dominant over $S$ such that each fiber $Z_{s}$ is a homologically trivial $n$-cocycle of $X_{s}$.
@@
(C.5) $\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{n}(X/S),\qquad s\mapsto{\mathrm{AJ}}(Z_{s})$
is an admissible normal function.
The following proposition is a simple application of the Abel–Jacobi map and the Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on intermediate Jacobians introduced in §$\S$§C.2. Recall the definition $S^{\mathrm{amp}}:=S\setminus S^{\mathrm{Betti}}(1)$.
@@
Assume $X/S$ and $Z$ are defined over $\overline{\mathbb{Q}}$. Then $[Z_{s}]\in\mathrm{Ch}^{n}(X_{s})$ is non-torsion for any $s\in S^{\mathrm{amp}}({\mathbb{C}})\setminus S^{\mathrm{amp}}(\overline{\mathbb{Q}})$.
@@
Assume $[Z_{s}]$ is a torsion point of $\mathrm{Ch}^{n}(X_{s})$. Then $[Z_{t}]$ is a torsion point of $\mathrm{Ch}^{n}(X_{t})$ for each $t\in\bar{s}(\overline{\mathbb{Q}})$, and $t\mapsto{\mathrm{AJ}}(Z_{t})$ is a torsion section of ${\mathscr{J}}^{n}(X/S)\times_{S}\bar{s}\rightarrow\bar{s}$. So $\nu_{Z}(\bar{s})\subseteq{\mathscr{F}}_{\mathrm{Betti}}$. If $s\not\in S(\overline{\mathbb{Q}})$, then $\dim\bar{s}\geq 1$, and hence $\bar{s}\subseteq S^{\mathrm{Betti}}(1)$ by definition. Now we are done. ∎
@@ C.6. Normal functions associated with Gross–Schoen cycles and Ceresa cycles
Let $C$ be an irreducible smooth projective curve defined over a field $k$ of genus $g\geq 3$. Let $e\in C(k)$. We have:
    For each subset $T\subseteq\{1,2,3\}$, the modified diagonal $\Delta_{T}(C)=\{(x_{1},x_{2},x_{3}):x_{i}=e\text{ for }i\not\in T,~{}x_{j}=x_{j^{\prime}}\text{ for all }j,j^{\prime}\in T\}\in\mathrm{Ch}^{2}(C^{3})$.
    The (classical) Abel–Jacobi map $i_{e}\colon C\rightarrow\mathrm{Jac}(C)$ sending $x\mapsto[x-e]$.
Both $\Delta_{T}(C)$ and $i_{e}$ can be extended to any $e=\sum_{i}n_{i}e_{i}\in\mathrm{Div}^{1}(C)(k)=\mathrm{Pic}^{1}(C)(k)$. This is classical and direct for $i_{e}$, and for $\Delta_{T}(C)$ one can define $\Delta_{123}(C)=\{(c,c,c):c\in C\}$, $\Delta_{12}(C)=\sum n_{i}\{(c,c,e_{i}):c\in C\}$, $\Delta_{1}(C)=\sum_{i,j}n_{i}n_{j}\{(c,e_{i},e_{j}):c\in C\}$, etc.
For any $e\in\mathrm{Pic}^{1}(C)$, define the Gross–Schoen cycle based at $e$ (resp. the Ceresa cycle based at $e$) to be:
  • (Gross–Schoen cycle based at $e$) $\Delta_{{\mathrm{GS}},e}(C):=\Delta_{123}(C)-\Delta_{12}(C)-\Delta_{13}(C)-\Delta_{23}(C)+\Delta_{1}(C)+\Delta_{2}(C)+\Delta_{3}(C)\in\mathrm{Ch}^{2}(C^{3})$.
  • (Ceresa cycle based at $e$) ${\mathrm{Ce}}_{e}(C):=i_{e}(C)-[-1]_{*}i_{e}(C)\in\mathrm{Ch}^{g-1}(\mathrm{Jac}(C))$.
${\mathrm{Ce}}_{e}(C)$ is clearly homologically trivial because $[-1]$ acts trivially on even-degree cohomology groups. It is not hard to check that $\Delta_{{\mathrm{GS}},e}(C)$ is also homologically trivial; this follows for example from [GS95, Prop. 3.1] and because the map $\mathrm{Pic}^{1}(C)\rightarrow\mathrm{Ch}^{2}(C^{3})$, $e\mapsto\Delta_{{\mathrm{GS}},e}(C)$ is a group homomorphism.
Let $\omega_{C}$ be the canonical divisor on $C$ and $[\omega_{C}]$ its divisor class.
@@
There are finitely many $\xi$’s with $(2g-2)\xi=[\omega_{C}]$ and each two differ from a $(2g-2)$-torsion. Hence, the Gross–Schoen cycle and the Ceresa cycle of $C$ are well-defined at $(2g-2)$-torsions.
We will associate normal functions with the Gross–Schoen cycles and Ceresa cycles. First we have ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}^{2}(C^{3})={\mathscr{J}}(H^{3}(C^{3},{\mathbb{Z}})(2))$ and ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}^{g-1}(\mathrm{Jac}(C))={\mathscr{J}}(H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1))$. Using the Poincaré duality on $\mathrm{Jac}(C)$ and on $C$, we get $H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1)=H_{3}(\mathrm{Jac}(C),{\mathbb{Z}})(-1)=\bigwedge^{3}H_{1}(C,{\mathbb{Z}})(-1)=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)$. So ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}(\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2))$.
We have $\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)\subseteq H^{3}(C^{3},{\mathbb{Z}})(2)$ in the following way. First, the Künneth formula gives a decomposition $H^{3}(C^{3},{\mathbb{Z}})=H^{1}(C,{\mathbb{Z}})^{\otimes 3}\bigoplus H^{1}(C,{\mathbb{Z}})(-1)^{\oplus 6}$. Next, $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$, as a subspace of $H^{3}(C^{3},{\mathbb{Z}})$, has a basis consisting of $\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}$, with $\alpha_{j}$ the pullback of an element in $H^{1}(C,{\mathbb{Z}})$ under the $j$-th projection $C^{3}\rightarrow C$. The symmetric group $S_{3}$ acts naturally on $C^{3}$, and this induces an action of $S_{3}$ on $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$ with
$\sigma(\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3})=\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$
for each $\sigma\in S_{3}$. Then $(H^{1}(C,{\mathbb{Z}})^{\otimes 3})^{S_{3}}=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})$ for this action, with each member having a basis consisting of $\sum_{\sigma\in S_{3}}\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$.
@@
For appropriate numbers $a_{j}$. The multiplication $[n]\colon\mathrm{Jac}(C)\rightarrow\mathrm{Jac}(C)$ induces
$\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})\rightarrow\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}}),\quad x\mapsto n^{3}x.$
Hence $n^{2+j}{\mathrm{AJ}}(C_{j})=n^{3}{\mathrm{AJ}}(C_{j})$ for all $n\in{\mathbb{Z}}$, and so ${\mathrm{AJ}}(C_{j})=0$ for all $j\geq 2$.
@@
Now, we turn to defining and studying the normal functions. Let ${\mathscr{M}}_{g}$ be the moduli space of curves of genus $g$, and let $f\colon{\mathscr{C}}_{g}\rightarrow{\mathscr{M}}_{g}$ be the universal curve. We have a VHS $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}\rightarrow{\mathscr{M}}_{g}$. The fiberwise polarization
$q_{H}\colon H^{1}(C,{\mathbb{Z}})\otimes H^{1}(C,{\mathbb{Z}})\xrightarrow{\cup}H^{2}(C,{\mathbb{Z}})\simeq{\mathbb{Z}}(-1),$
with $\cup$ the cup product (the dual of the intersection pairing) on $H^{1}(C,{\mathbb{Z}})$, induces a polarization on $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}$.